# Exploring Feature-based Knowledge Distillation For Recommender System: A Frequency Perspective

### v1

@@ Abstract.
In this paper, we analyze the feature-based knowledge distillation for recommendation from the frequency perspective. By defining knowledge as different frequency components of the features, we theoretically demonstrate that regular feature-based knowledge distillation is equivalent to equally minimizing losses on all knowledge and further analyze how this equal loss weight allocation method leads to important knowledge being overlooked. In light of this, we propose to emphasize important knowledge by redistributing knowledge weights. Furthermore, we propose FreqD, a lightweight knowledge reweighting method, to avoid the computational cost of calculating losses on each knowledge. Extensive experiments demonstrate that FreqD consistently and significantly outperforms state-of-the-art knowledge distillation methods for recommender systems. Our code is available at https://anonymous.4open.science/r/FreqKD/.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Collaborative filtering
@@ 1. Introduction
With the rapid growth in the variety of items and the gradual diversification of user preferences, recent studies (Ohsaka and Togashi, 2023; Liang et al., 2018) have begun to increase the model capacity by increasing the dimensionality of user/item embeddings. However, due to the huge share of embeddings (Ohsaka and Togashi, 2023; Zhao et al., 2023) in the overall model parameters, increasing the embedding dimensionality greatly increases the number of parameters, which in turn increases the storage cost and inference latency, leading to longer waiting times and lower user satisfaction.
@@ 2.3. Knowledge Distillation in Recommender System
Kowledge distillation has been introduced into recommender systems to reduce the representation dimensionality and inference latency, which falls into the same three categories, i.e., response-based, feature-based, and relation-based.
Response-based methods focus on teachers’ predictions. For example, CD (Lee et al., 2019) samples unobserved items from a distribution associated with their rankings predicted by students: items with higher rankings are more likely to be sampled. RRD (Kang et al., 2020) adopts a list-wise loss to maximize the likelihood of teachers’ recommendation list. UnKD (Chen et al., 2023) partitions items into multiple groups based on popularity and samples positive-negative item pairs within each group.
@@
Relation-based methods focus on the relationships between different entities (i.e., users and items). HTD (Kang et al., 2021) observes that the student is too small to learn the whole relationship. As a result, the vanilla relation-based distillation approach is not always effective and even degrades the student’s performance. Therefore, it proposes to distill the sample relation hierarchically to alleviate the large capacity gap between the student and the teacher. Since relation-based knowledge distillation methods are also closely related to features, we discuss how the main idea of this paper could be applied to them in Appendix A.5.
@@ 3.2. Feature-based Knowledge Distillation
Let $d^{S}$ and $d^{T}$ denote the dimensionalities of student and teacher features, respectively. In this work, we use $\mathbf{S}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{I}|)\times d^{S}}$ and $\mathbf{T}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{I}|)\times d^{T}}$ to represent all the features of the student model and the teacher model, respectively.
@@
where $Proj(\cdot)$ denotes the projector that aligns the dimensionalities of $\mathbf{S}$ and $\mathbf{T}$, and is usually a linear layer.
3.3. Graph Signal Processing
In the following section, we perform the frequency analysis on the user-item bipartite graph constructed from observed user-item interactions. Thus, the number of nodes is equal to the sum of the number of users and the number of items, i.e., $|\mathcal{V}|=|\mathcal{U}|+|\mathcal{I}|$.
@@ 4. Methodology
In Section 4.1, we define the knowledge in features and show the difference between various types of knowledge. In Section 4.2, we propose a simple yet effective method to reweight the distillation loss of different knowledge. In Section 4.3, we further propose FreqD, a lightweight knowledge reweight method. The overall training process of FreqD is exhibited in Algorithm 1 and the time complexity analysis is provided in Appendix A.3. A further discussion on extending our analytical approach to relation-based knowledge distillation methods is provided in Appendix A.5.
@@ 5.1. Experimental Settings
Implementation Details. We implement all the methods with PyTorch and use Adam as the optimizer in all experiments. For our method, the weight decay is selected from {1e-3, 1e-4, 1e-5, 0}. The search space of the learning rate is {1e-3, 1e-4}. $\beta$ is selected from {0.01, 0.05, 0.1, 0.5}. We set the total number of training epochs as 1000. We also conduct early stopping according to the NDCG@20 on the validation set and stop training when the NDCG@20 doesn’t increase for 30 consecutive epochs. All hyperparameters of the compared baselines are tuned to ensure optimal performance.
(10) $h(\lambda)=1-\alpha\lambda,$
where $0\leq\alpha\leq 0.5$ is a hyperparameter. We have found it to be well-behaved while being very simple and efficient. Moreover, in Section 5.5, we provide the performance when $h(\lambda)$ is set to a quadratic function form for further verification. Some key hyperparameter analysis is provided in Appendix A.7.
@@
5.2. Performance Comparison
Overall Performance. The results shown in Table 4 indicate that:
Ablation Study. Since FitNet can be regarded as a special case of our method obtained by equally weighting losses on all types of knowledge, here we serve as an ablation experiment by comparing our method with FitNet. By comparing the performance of FitNet and FreqD in Table 4, we find that although FreqD is simple, it significantly outperforms FitNet on all datasets and all backbones. This suggests that by decoupling the knowledge contained in the features and controlling the weights of each type of knowledge separately, we can effectively facilitate the learning of more important knowledge by the student model and consequently gain improved results. It also shows that the conclusions we get from decoupling the various types of knowledge in features and analyzing their performance in feature-based distillation are useful for designing practical methods.
@@ 5.4. Inference Efficiency
From the results in Table 7, it is obvious that our method can benefit students of all sizes. Moreover, we achieve comparable performance to the teacher with only 10% of the number of parameters and about a quarter of the inference latency. When the student is half the size of the teacher, the student can significantly outperform the teacher with only half the storage cost and greatly increase the inference efficiency. Moreover, with the same number of parameters, our method significantly outperforms the teacher, demonstrating that FreqD is also effective in improving existing recommendation models.
@@ 5.5. Different Choices of $h(\lambda)$
In Eq.(10), we set $h(\lambda)$ as a linear function for simplicity and efficiency. However, it is possible to design it into other higher-order forms. In this section, we verify empirically that the linear form, despite its simplicity, has a good performance. In Table 8, we give the results when $h(\lambda)$ is designed as a quadratic function, i.e., $h(\lambda)=a\lambda^{2}+b\lambda+1$. Note that we similarly restrict its shape to make it decreasing in the domain of definition. From the results, we find that modifying $h(\lambda)$ to a quadratic function can slightly improve recommendation accuracy on Gowalla, while on Yelp it performs similarly to when it is a linear function. Therefore, we believe that the linear function of $h(\lambda)$ can be used as a strong option.
@@ Appendix A Appendix
A.1. Proof of Theorem 2
@@ A.2. Proof of Theorem 3
A.3. Time Complexity
A.4. Training Process
Input : Training data $\mathcal{R}$; Pretrained teacher model $T$.
Output : Student model $S$.
3 Compute the graph filter $\mathcal{H}(\tilde{\mathbf{L}})$ using Eq.(7).
4 while not convergence do
5       for each batch $\mathcal{B}\in\mathcal{R}$ do
6             Compute $\mathcal{L}_{Base}$.
9             Compute $\mathcal{L}_{FreqD}$ according to Eq.(8).
10             Compute $\mathcal{L}=\mathcal{L}_{Base}+\beta\cdot\mathcal{L}_{FreqD}$.
11             Update the Student model and the projector by minimizing $\mathcal{L}$.
12       end for
13
14 end while
Algorithm 1 FreqD
A.5. Further Discussion
Theorem 1.
Given the widely used relation-based distillation loss
$\mathcal{L}_{RD}=\|\mathbf{SS}^{\top}-\mathbf{TT}^{\top}\|_{F}^{2}$
A.6. Proof of Theorem 1

### v2

@@ Abstract.
In this paper, we analyze the feature-based knowledge distillation for recommendation from the frequency perspective. By defining knowledge as different frequency components of the features, we theoretically demonstrate that regular feature-based knowledge distillation is equivalent to equally minimizing losses on all knowledge and further analyze how this equal loss weight allocation method leads to important knowledge being overlooked. In light of this, we propose to emphasize important knowledge by redistributing knowledge weights. Furthermore, we propose FreqD, a lightweight knowledge reweighting method, to avoid the computational cost of calculating losses on each knowledge. Extensive experiments demonstrate that FreqD consistently and significantly outperforms state-of-the-art knowledge distillation methods for recommender systems. Our code is available at https://github.com/woriazzc/KDs.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709248isbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Information systems Retrieval efficiency
@@ 1. Introduction
With the rapid growth in the variety of items and the gradual diversification of user preferences, recent studies (Ohsaka and Togashi, 2023; Liang et al., 2018) have begun to increase the model capacity by increasing the dimension of user and item embeddings. However, due to the huge share of embeddings (Ohsaka and Togashi, 2023; Zhao et al., 2023) in the overall model parameters, increasing the embedding dimensions greatly increases the number of parameters, which in turn increases the storage cost and inference latency, leading to longer waiting times and lower user satisfaction.
@@ 2.3. Knowledge Distillation in Recommender System
Knowledge distillation has been introduced into recommender systems to reduce the representation dimensionality and inference latency, which falls into the same three categories, i.e., response-based, feature-based, and relation-based.
Response-based methods focus on teachers’ predictions. For example, CD (Lee et al., 2019) samples unobserved items from a distribution associated with their rankings predicted by students: items with higher rankings are more likely to be sampled. RRD (Kang et al., 2020) adopts a list-wise loss to maximize the likelihood of teachers’ recommendation list. DCD (Lee and Kim, 2021) samples items and users from distributions associated with teacher-student prediction discrepancy to provide the student with dynamic knowledge. UnKD (Chen et al., 2023) partitions items into multiple groups based on popularity and samples positive-negative item pairs within each group. HetComp (Kang et al., 2023) aims to transfer the ensemble knowledge of heterogeneous teachers. It guides the student model by transferring easy-to-hard knowledge sequences generated from the teachers’ trajectories. CCD (Lee et al., 2024) focuses on KD in a non-stationary data stream, where both the teacher and the student continually and collaboratively evolve along the data stream. It makes the student and teacher promote each other.
@@
Relation-based methods focus on the relationships between different entities (i.e., users and items). HTD (Kang et al., 2021) observes that the student is too small to learn the whole relationship. As a result, the vanilla relation-based distillation approach is not always effective and even degrades the student’s performance. Therefore, it proposes to distill the sample relation hierarchically to alleviate the large capacity gap between the student and the teacher. Since relation-based knowledge distillation methods are also closely related to features, we discuss how the main idea of this paper could be applied to them in Section 4.5.
@@ 3.1. Top-N Recommendation
3.2. Graph Signal Processing
In the following section, we perform the frequency analysis on the user-user KNN graph and the item-item KNN graph separately. We construct them from the predictions of the teacher model. Specifically, before training, we calculate the Euclidean distances between each pair of users (or items) using the frozen teacher model and obtain the KNN graph. Thus, when performing our method on the user KNN graph (or the item KNN graph), the number of nodes is equal to the number of users (or items, respectively).
@@ 3.3. Feature-based Knowledge Distillation
Let $d^{S}$ and $d^{T}$ denote the dimensions of student and teacher features, respectively. In this work, we use $\mathbf{S}\in\mathbb{R}^{|\mathcal{V}|\times d^{S}}$ and $\mathbf{T}\in\mathbb{R}^{|\mathcal{V}|\times d^{T}}$ to represent all the features of the student model and the teacher model, respectively. Note that, when performing knowledge for users, we have $|\mathcal{V}|=|\mathcal{U}|$, otherwise $|\mathcal{V}|=|\mathcal{I}|$.
@@
where $Proj(\cdot)$ denotes the projector that aligns the dimensionalities of $\mathbf{S}$ and $\mathbf{T}$, and is usually a linear transformation.
@@ 4. Methodology
In Section 4.1, we define the knowledge in features and show the difference between various types of knowledge. In Section 4.2, we propose a simple yet effective method to reweight the distillation loss of different knowledge. In Section 4.3, we further propose FreqD, a lightweight knowledge reweight method. The overall training process of FreqD is exhibited in Algorithm 1 and the time complexity analysis is provided in Section 4.4. We further discuss the relationship of our method with FitNet in Section 4.5. Finally, we also discuss the prospect of extending our analytical approach to relation-based knowledge distillation methods in Appendix A.1.
@@ 4.3. Lightweight Knowledge Reweighting Method
Input : Training data $\mathcal{R}$; Pretrained teacher model $T$.
Output : Student model $S$.
3 Compute the graph filter $\mathcal{H}(\tilde{\mathbf{L}})$ using Eq.(7).
4 while not convergence do
5       for each batch $\mathcal{B}\in\mathcal{R}$ do
6             Compute $\mathcal{L}_{Base}$.
9             Compute $\mathcal{L}_{FreqD}$ according to Eq.(8).
10             Compute $\mathcal{L}=\mathcal{L}_{Base}+\beta\cdot\mathcal{L}_{FreqD}$.
11             Update the Student model and the projector by minimizing $\mathcal{L}$.
12       end for
13
14 end while
Algorithm 1 FreqD
4.4. Time Complexity
4.5. Further Discussion
Relation to FitNet. FitNet (Romero et al., 2014) is a widely used method for feature-based KD. It simply uses conventional feature-based knowledge distillation loss. It can regarded be as a special case of our method. Specifically, we can rewrite Eq.(8) as follows:
(10) $\mathcal{L}_{FreqD}$ $=\sum_{k=1}^{|\mathcal{V}|}h^{2}(\lambda_{k})\|(Proj(\tilde{\mathbf{S}}_{k})-\tilde{\mathbf{T}}_{k})\|_{F}^{2}$
(11) $=\left\|\left(\sum_{k=0}^{K}\theta_{k}\tilde{\mathbf{L}}^{k}\right)Proj(\textbf{S})-\left(\sum_{k=0}^{K}\theta_{k}\tilde{\mathbf{L}}^{k}\right)\textbf{T}\right\|_{F}^{2}.$
It follows that FitNet can be regarded as a 0-order special case of our approach. Specifically, by setting the hyperparameter $K$ to 00, we can exactly recover FitNet. It also means that all frequency components are weighted equally in FitNet, i.e., $h^{2}(\lambda_{k}),k=1,\cdots,|\mathcal{V}|$ is constant. Combined with our previous analysis, it can be assumed that we have improved FitNet based on the fact that “low frequency” is more important in the field of recommender systems.
@@ 5. Experiments
In this section, we conduct experiments on three public datasets using three backbones to validate the effectiveness of FreqD. The overall performance comparison and ablation study are presented in Section 5.2. The analysis of training efficiency and inference efficiency are presented in Section 5.3 and Section 5.4, respectively. In Section 5.5, we provide the performance when $h(\lambda)$ is set to a quadratic function form. Some key hyperparameter analysis is provided in Appendix A.5.
@@ 5.1. Experimental Settings
Implementation Details. We implement all the methods with PyTorch and use Adam as the optimizer in all experiments. For our method, the weight decay is selected from {1e-3, 1e-4, 1e-5, 0}. The search space of the learning rate is {1e-3, 1e-4}. $\beta$ is selected from {0.01, 0.05, 0.1, 0.5}. We set the total number of training epochs as 1000. We also conduct early stopping according to the NDCG@20 on the validation set and stop training when the NDCG@20 doesn’t increase for 30 consecutive epochs. All hyperparameters of the compared baselines are tuned to ensure optimal performance. Before graph filtering on the KNN graph, we perform edge dropout, following the practice in LightGCN (He et al., 2020) to alleviate overfitting.
(12) $h(\lambda)=1-\alpha\lambda,$
where $0\leq\alpha\leq 0.5$ is a hyperparameter. We have found it to be well-behaved while being very simple and efficient. Moreover, in Section 5.5, we provide the performance when $h(\lambda)$ is set to a quadratic function form for further verification.
@@
5.2. Performance Comparison
Overall Performance. The results shown in Table 4 indicate that:
Ablation Study. Since FitNet can be regarded as a special case of our method obtained by equally weighting losses on all types of knowledge, here we serve as an ablation experiment by comparing our method with FitNet. By comparing the performance of FitNet and FreqD in Table 4, we find that although FreqD is simple, it significantly outperforms FitNet on all datasets and all backbones. This suggests that by decoupling the knowledge contained in the features and controlling the weights of each type of knowledge separately, we can effectively facilitate the learning of more important knowledge by the student model and consequently gain improved results. It also shows that the conclusions we get from decoupling the various types of knowledge in features and analyzing their performance in feature-based distillation are useful for designing practical methods. We believe that our design concept will be useful for other knowledge distillation methods in the field of recommender systems.
@@ 5.4. Inference Efficiency
From the results in Table 7, it is obvious that our method can benefit students of all sizes. Moreover, we achieve comparable performance to the teacher with only 10% of the number of parameters and about a quarter of the inference latency. It demonstrates that by distilling with our method, we can significantly improve the inference efficiency of existing recommendation models. When the student is half the size of the teacher, the student can significantly outperform the teacher with only half the storage cost and greatly increase the inference efficiency. Moreover, with the same number of parameters, our method significantly outperforms the teacher, demonstrating that FreqD is also effective in improving existing recommendation models.
@@ 5.5. Different Choices of $h(\lambda)$
In Eq.(12), we set $h(\lambda)$ as a linear function for simplicity and efficiency. However, it is possible to design it into other higher-order forms. In this section, we verify empirically that the linear form, despite its simplicity, has a good performance. In Table 8, we give the results when $h(\lambda)$ is designed as a quadratic function, i.e., $h(\lambda)=a\lambda^{2}+b\lambda+1$. Note that we similarly restrict its shape to make it decrease in the domain of definition. From the results, we find that modifying $h(\lambda)$ to a quadratic function can slightly improve recommendation accuracy on Gowalla, while on Yelp it performs similarly to when it is a linear function. Therefore, we believe that the linear function of $h(\lambda)$ can be used as a strong option.
@@ 7. Conclusion
8. Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 92270119, Grant 62072182, and the Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education.
@@ Appendix A Appendix
A.1. Prospects for applications on Relation-based KD
Theorem 1.
Given the widely used relation-based distillation loss
(13) $\mathcal{L}_{RD}=\|\mathbf{SS}^{\top}-\mathbf{TT}^{\top}\|_{F}^{2}$
A.2. Proof of Theorem 2
@@ A.3. Proof of Theorem 3
A.4. Proof of Theorem 1